# Representability of the direct sum of uniform $q$-matroids

### v1

@@ 4.2. Construction for two summands over small extension fields
Let $p$ be the characteristic of $\mathbb{F}_{q^{m}}$, $m=p^{r}$, and let $n_{1},n_{2}$ be any positive integers such that $n_{1}+n_{2}-1\leq\frac{m}{2}$. Furthermore, let $\xi\in\mathbb{F}_{q^{m}}^{*}$ be such that $\mathcal{F}_{n_{1}+n_{2}-1}\cap\xi\mathcal{F}_{n_{1}+n_{2}-1}=\{0\}$. Define $\mathcal{S}_{1}:=\mathcal{F}_{n_{1}}$ and $\mathcal{S}_{2}:=\{a+\xi a^{q}\colon a\in\mathcal{F}_{n_{2}}\}$. Then, $\mathcal{S}_{1}\oplus\mathcal{S}_{2}$ is a $(\Lambda_{1,(1,1)},1)$-evasive $[n_{1}+n_{2},2]_{q^{m}/q}$ system.
@@
Combining Theorem 2.4 with Proposition 4.6, [25, Theorem 4.5], Corollary 4.8 and Corollary 4.11, we obtain the $\mathbb{F}_{q^{m}}$-representability of the direct sum $\mathcal{U}_{1,n_{1}}(q)\oplus\mathcal{U}_{2,n_{2}}(q)$ under the assumptions (1)-(5) of the statement. ∎
@@
Let $n_{2}=2$ and let $n_{1}$ be any positive integer greater than $1$. Combining Corollary 4.2 and Theorem 4.4(2), we obtain that $\mathcal{U}_{1,n_{1}}(q)\oplus\mathcal{U}_{1,2}(q)$ is $\mathbb{F}_{q^{m}}$-representable if and only if $m\geq 2n_{1}$. This generalizes the characterization result for the $\mathbb{F}_{q^{m}}$-representability of two uniform $q$-matroids of rank $k_{1}=k_{2}=1$ and height $n_{1}=n_{2}=2$ given [17, Proposition 3.8] (see Example 1.16).
@@
The smallest case left out by Theorem 4.4 is whether the direct sum of two uniform $q$-matroids of rank $1$ and height $3$ is $\mathbb{F}_{q^{7}}$-representable. Using the algebra software magma [7], we found examples of $(\Lambda_{1,(1,1)},1)$-evasive subspaces for $q\in\{2,3,4,5,7\}$. Theorem 2.4 implies that the direct sum of two copies of $\mathcal{U}_{1,3}(q)$ is $\mathbb{F}_{q^{7}}$-representable for $q\in\{2,3,4,5,7\}$.
@@
Let $n_{1}=7$ and $n_{2}=6$. We know by Corollary 4.2 that $\mathcal{M}=\mathcal{U}_{1,7}(q)\oplus\mathcal{U}_{1,6}(q)$ is not $\mathbb{F}_{q^{m}}$-representable for each $m\in[14]$. On the other hand, $\mathcal{M}$ is $\mathbb{F}_{q^{m}}$-representable for every $m\geq 42$ (Theorem 4.4(2)) and for $m\in\{16,18,20,\ldots,38,40\}$ (Theorem 4.4(1)). Moreover, using Theorem 4.4(4), we also derive that $\mathcal{M}$ is $\mathbb{F}_{q^{m}}$-representable for $m\in\{21,27,33,35,39\}$. Finally, Theorem 4.4(5) implies that $\mathcal{M}$ is $\mathbb{F}_{q^{m}}$-representable also when
$m=\begin{cases}25&\mbox{ if }q=5^{h},\\ 29&\mbox{ if }q=29^{h},\\ 31&\mbox{ if }q=31^{h},\\ 37&\mbox{ if }q=37^{h},\\ 41&\mbox{ if }q=41^{h}.\end{cases}$ (17)
The only open cases are for $m\in\{17,19,23,25^{*},29^{*},31^{*},37^{*},41^{*}\}$, where the indicates that the case is open except for some values of the characteristic of the field, according to (17).

### v2

@@ 4.2. Construction for two summands over small extension fields
Let $p$ be the characteristic of $\mathbb{F}_{q^{m}}$, $m=p^{r}$, and let $n_{1},n_{2}$ be any positive integers such that $n_{1}+n_{2}-1\leq\frac{m}{2}$. Furthermore, let $\xi\in\mathbb{F}_{q^{m}}^{*}$ be such that $\mathcal{F}_{n_{1}+n_{2}-1}\cap\xi\mathcal{F}_{n_{1}+n_{2}-1}=\{0\}$. Define $\mathcal{S}_{1}:=\mathcal{F}_{n_{1}}$ and $\mathcal{S}_{2}:=\{a+\xi a^{q}\colon a\in\mathcal{F}_{n_{2}}\}$. Then, $\mathcal{S}_{1}\oplus\mathcal{S}_{2}$ is a $(\Lambda_{1,(1,1)},1)$-evasive $n_{1}+n_{2},2]_{q^{m}/q}$ system.
@@
By Theorem 2.4, the direct sum $\mathcal{U}_{1,n_{1}}(q)\oplus\mathcal{U}_{1,n_{2}}(q)$ is $\mathbb{F}_{q^{m}}$-representable if and only if there exist $\mathbb{F}_{q}$-subspaces $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$ of $\mathbb{F}_{q^{m}}$ such that $\mathcal{S}_{1}\oplus\mathcal{S}_{2}$ is $(\Lambda_{1,(1,1)},1)$-evasive. The existence of such subspaces is ensured:
  • by Corollary 4.8, under assumptions (1) and (4),
  • by Proposition 4.6, under assumption (2),
  • by [25, Theorem 4.5], under assumption (3), and
  • by Corollary 4.11, under assumption (5).
@@
Let $n_{2}=2$ and let $n_{1}$ be any positive integer greater than $1$. Combining Corollary 4.2 and Theorem 4.4(2), we obtain that $\mathcal{U}_{1,n_{1}}(q)\oplus\mathcal{U}_{1,2}(q)$ is $\mathbb{F}_{q^{m}}$-representable if and only if $m\geq 2n_{1}$. This generalizes the characterization result for the $\mathbb{F}_{q^{m}}$-representability of two uniform $q$-matroids of rank $k_{1}=k_{2}=1$ and height $n_{1}=n_{2}=2$ given [19, Proposition 3.8] (see Example 1.16).
@@
The smallest case left out by Theorem 4.4 is whether the direct sum of two uniform $q$-matroids of rank $1$ and height 33 is $\mathbb{F}_{q^{7}}$-representable. Using the algebra software magma [7], we found examples of $(\Lambda_{1,(1,1)},1)$-evasive subspaces for $q\in\{2,3,4,5,7\}$. Theorem 2.4 implies that the direct sum of two copies of $\mathcal{U}_{1,3}(q)$ is $\mathbb{F}_{q^{7}}$-representable for $q\in\{2,3,4,5,7\}$.
@@
Let $n_{1}=7$ and $n_{2}=6$. We know by Corollary 4.2 that $\mathcal{M}=\mathcal{U}_{1,7}(q)\oplus\mathcal{U}_{1,6}(q)$ is not $\mathbb{F}_{q^{m}}$-representable for each $m\in[12]$. On the other hand, $\mathcal{M}$ is $\mathbb{F}_{q^{m}}$-representable for every $m\geq 42$ (Theorem 4.4(2)) and for $m\in\{14,16,18,20,\ldots,38,40\}$ (Theorem 4.4(1)). Moreover, using Theorem 4.4(4), we also derive that $\mathcal{M}$ is $\mathbb{F}_{q^{m}}$-representable for $m\in\{21,27,33,35,39\}$. Finally, Theorem 4.4(5) implies that $\mathcal{M}$ is $\mathbb{F}_{q^{m}}$-representable also when
[…]
The only open cases are for $m\in\{13,15,17,19,23,25^{*},29^{*},31^{*},37^{*},41^{*}\}$, where the indicates that the case is open except for some values of the characteristic of the field, according to (19).